Outer position loop D(s) Inner attitude loo V(s) 10(s) x,(s) 의 1 X(s) jar(s) 1、 1

Assuming the closed-loop system is stable, find the steady-state error if

0 0
Answer #1

Inner altitude Loop:

There is another loop inside this:

Forward gain of this loop is:

G(s) = \frac{1}{Js}

Feedback gain:

H(s) = Ku

closed loop transfer function:

P()+GH

P(s) = \frac{\frac{1}{Js}K_w}{1+\frac{1}{Js}K_w}

P(s) =

Now consider the whole inner altitude loop:

Forward gain of this loop is:

P(s)*\frac{1}{s}*k_{\theta} = \frac{K_w}{Js+K_w}*\frac{1}{s}*k_{\theta}

Feedback gain is 1

closed loop transfer function:

Q(s) =\frac{\theta(s)}{\theta_r(s)}= \frac{ \frac{K_w}{Js+K_w}*\frac{1}{s}*k_{\theta}}{1+ \frac{K_w}{Js+K_w}*\frac{1}{s}*k_{\theta}}

θ(s)

0(s)

Now consider the loop on the right side:

Forward gain of this loop is:

F(s) =

Feedback gain is

B(s) = b

closed loop transfer function:

F B

E(s) = \frac{\frac{1}{s}*b}{1+\frac{1}{s}*b}

E(s) = \frac{b}{s+b}

Now the block diagram can be reduced as:

D(s) X(s) Q(s)

Let the whole part circled in blue be:

Tis

\small X(s) = ((X_r-X(s))*I(s)+D(s))E(s)*(1/s)

But if, X_r(s) = 0,

\small X(s) = (-X(s)*I(s)+D(s))E(s)*(1/s)

\small X(s) = -X(s)*I(s)*E(s)*(1/s)+D(s)*E(s)*(1/s)

\small X(s)(1+I(s)*E(s)*(1/s))=D(s)*E(s)*(1/s)

\small \Rightarrow X(s)=\frac{D(s)*E(s)}{s+I(s)*E(s)}

E(s) = \frac{b}{s+b}  

Disturbance is a step function: D(s) = \frac{1}{s}

\small \Rightarrow X(s)=\frac{\frac{1}{s}*\frac{b}{s+b}}{s+I(s)*\frac{b}{s+b}}

Tis

Q(s) = \frac{K_wk_{\theta}}{Js^2+K_ws+ K_wk_{\theta}}

\small I(s)=K_p(1+T_ds+\frac{1}{T_is})* \frac{K_wk_{\theta}}{Js^2+K_ws+ K_wk_{\theta}}*g

\small I(s)=K_p(\frac{T_is+T_iT_ds^2+1}{T_is})* \frac{K_wk_{\theta}}{Js^2+K_ws+ K_wk_{\theta}}*g

I(s) ak

\small \Rightarrow X(s)=\frac{\frac{1}{s}*\frac{b}{s+b}}{s+I(s)*\frac{b}{s+b}}

\small \Rightarrow X(s)=\frac{\frac{1}{s}*\frac{b}{s+b}}{s+\frac{(T_is+T_iT_ds^2+1)K_pK_wk_{\theta}}{(T_is)(Js^2+K_ws+ K_wk_{\theta)}}*g*\frac{b}{s+b}}

\small \Rightarrow X(s)=\frac{\frac{1}{s}b*(T_is)(Js^2+K_ws+ K_wk_{\theta})}{s*(s+b)(T_is)(Js^2+K_ws+ K_wk_{\theta})+(T_is+T_i T_d s^2+1)K_pK_wk_{\theta}*g*b}

By final value theorem:

\small X_{ss} = \lim_{s\rightarrow 0}sX(s)

\small \Rightarrow X_{ss}=\lim_{s\rightarrow 0}s*\frac{\frac{1}{s}b*(T_is)(Js^2+K_ws+ K_wk_{\theta})}{s*(s+b)(T_is)(Js^2+K_ws+ K_wk_{\theta})+(T_is+T_i T_d s^2+1)K_pK_wk_{\theta}*g*b}

\small \Rightarrow X_{ss}=\lim_{s\rightarrow 0}\frac{b*(T_is)(Js^2+K_ws+ K_wk_{\theta})}{s*(s+b)(T_is)(Js^2+K_ws+ K_wk_{\theta})+(T_is+T_i T_d s^2+1)K_pK_wk_{\theta}*g*b}

\small \Rightarrow X_{ss} = 0